# LLM2CLIP: Powerful Language Model Unlocks Richer Visual Representation

### v3

@@ Abstract
CLIP is a foundational multimodal model that aligns image and text features into a shared space using contrastive learning on large-scale image-text pairs. Its strength lies in leveraging natural language as a rich supervisory signal. With the rapid progress of large language models (LLMs), we explore their potential to further enhance CLIP’s multimodal representation learning. This work introduces a fine-tuning approach that integrates LLMs with the pretrained CLIP visual encoder, leveraging LLMs’ advanced text understanding and open-world knowledge to improve CLIP’s ability to process long and complex captions. To address the challenge of LLMs’ autoregressive nature, we propose a caption-to-caption contrastive learning framework to enhance the discriminative power of their outputs. Our method achieves substantial performance gains on various downstream tasks, demonstrating the effectiveness of combining LLMs with CLIP for enhanced multimodal learning.
@@ 1 Introduction
CLIP [32] is one of the most influential multimodal foundational models today. It aligns visual and language signals into a shared feature space by employing a simple contrastive learning loss on large-scale image-text pairs. As a retriever, CLIP supports a wide range of tasks, including zero-shot classification [31], detection [22], segmentation [46], and image-text retrieval [27, 17]. As a feature extractor, it has become dominant in virtually all cross-modal representation tasks, such as image understanding, video understanding, and text-to-image/video generation. For instance, works like LLaVA [24] and Qwen-VL [2] leverage CLIP as a feature extractor to obtain visual features for text models, while models like Stable Diffusion [34] and DALL·E 2 [33] use CLIP’s text encoder to extract textual features for visual models.
Figure 1: Zero-shot multimodal retrieval radar chart. LLM2CLIP significantly improves the performance of original models like CLIP and EVA02 through fine-tuning, substantially outperforming other fine-tuning-based methods such as LongCLIP.
@@
What makes CLIP particularly powerful as a vision encoder? The core of its strength lies in its unprecedented ability to align visual pretraining with natural language—the carrier of human knowledge. Unlike earlier vision encoders such as VGG and ResNet, which relied on the limited ImageNet dataset and simple image categories with just a few words, CLIP is trained on web-scale data using rich descriptive text. This alignment with language is what sets CLIP apart and unlocks its vast potential. However, since CLIP’s introduction, large language models (LLMs) have advanced significantly. Models like GPT-4 [1] and Llama [8] now demonstrate remarkable language capabilities, yet these advancements have not translated into corresponding improvements in visual representation learning. This raises the question: Can the capabilities of LLMs be harnessed to further improve multimodal representation learning?
@@
In this work, we aim to design a method that incorporates the powerful capabilities of LLMs to fine-tune pretrained state-of-the-art (SOTA) CLIP models, further enhancing their performance. We propose replacing the weak text encoder of CLIP with an LLM, followed by cross-modal contrastive learning training on image-text datasets. While LLMs have already been introduced into CLIP in the InternVL [6] framework, this approach requires large-scale retraining, and the actual performance improvement solely attributable to the LLM remains unclear. In our experiments, we found that directly using LLMs to fine-tune CLIP led to failure in all tasks, with a significant drop in performance across various downstream tasks as in Table 6. We discovered that the output features of LLMs for different texts (or even the same text) were very similar, which led to poor differentiation. This low discriminative power interferes with CLIP’s contrastive learning task. To investigate this further, we designed a caption-to-caption retrieval experiment, as shown in Table 1 and Figure 4. In this experiment, each image in the MS-COCO dataset has five human-annotated captions. We selected the first two captions as positive samples and performed retrieval across the entire test set. Using caption retrieval accuracy (CRA), we evaluated the text model’s ability to differentiate between captions, helping us determine which language model is better suited for CLIP. We found that Llama-3 8B achieved only 5.2% top-1 accuracy, while the standard CLIP-ViT-L reached 25.2% top-1 accuracy. As illustrated in Figure 4, the top-1 caption retrieved by the original Llama-3 model can be entirely unrelated to the query caption, clearly obstructing effective CLIP learning. Therefore, directly using an LLM to guide CLIP’s visual encoder training is highly constrained. Recent work on LLM2Vec also validates this idea, leading us to realize the necessity of fine-tuning the output space of LLMs.
Language Model CRA
@@
Table 1: Comparison of top-1 Caption Retrieval Accuracy (CRA) for various language models on the MS COCO 5K test set.
@@
We designed a caption contractive (CC) fine-tuning strategy, as Figure 2, pplying lightweight fine-tuning to the output tokens of Llama-3 8B using LoRA on the CC3M [36] image captioning dataset. The primary goal of this training task was to adjust the output space, improving the model’s ability to distinguish between different captions. We utilized a supervised SimCSE [11, 3] contrastive learning loss, where the original captions and re-annotated captions generated by ShareCaptioner [5] were treated as positive pairs, pulling them closer. In contrast, all other captions formed a negative sample set that the model learned to push away. Remarkably, after this CC fine-tuning, the caption retrieval accuracy, as shown in Table 1, rose from 5.2% to 29.5%. This successful fine-tuning process enables us to more effectively harness the capabilities of LLMs for CLIP training.
Building on fine-tuning of LLMs, we designed an efficient approach, LLM2CLIP, to enhance the performance of pretrained CLIP. As in  Figure 3, This method uses an LLM to further fine-tune the pretrained CLIP, allowing the LLM to impart knowledge and capabilities to CLIP’s vision encoder that were lacking during its initial pretraining. LLM2CLIP consists of two stages: in the first stage, we apply caption contrastive (CC) fine-tuning on the LLM with a small dataset to adjust the LLM’s output space. In the second stage, we fine-tune the CLIP vision encoder with the CC-fine-tuned LLM, with LLM gradients frozen, to further improve CLIP’s performance. Experimental results demonstrate that training on only 15M open-source data substantially boosts CLIP’s capabilities, with even 3M data yielding impressive results.
@@
    We propose LLM2CLIP, an efficient and effective LLM-assisted fine-tuning method for CLIP, enabling low-cost fine-tuning of pretrained CLIP models and achieving significant performance improvements.
@@ 2 Related Works
Several works have explored the integration of LLMs into CLIP. JinaCLIP [17] employed Jina-embeddings-v2 [13] as the text encoder, which is a BERT variant with 137M parameters, supporting longer texts. Though achieving similar visual performance to EVA-CLIP [38], its text encoder is far behind our used Llama3-8B, limiting the potential benefits from LLMs. E5-V [15] took a different approach by aggregating LLM features at the MLLM layer of LLAVA-next-8B [19] while freezing the ViT gradient to focus on adjusting the MLLM output. However, freezing visual encoder does not address the problem that ViT inherently lack complex visual feature extraction capabilities, leading to much worse performance than LLM2CLIP. MATE [14] designed a learnable adaptor to bridge the gap between CLIP’s text encoder and LLMs. They trained the CLIP visual encoder using LoRA on a small dataset focused on long-text image retrieval tasks. However, they did not recognize the critical issue we propose in this paper: the poor separability of LLM feature space, which is insufficient for direct support of CLIP training. Our work aims to thoroughly explore the impact of LLMs on CLIP as a foundational model, going beyond the influence on long-text retrieval.
CLIP meets Longer Captions.
@@
It has been widely recognized that the quality of CLIP’s text embedding is coarse and limited to only 77 tokens. Many works have attempted to extend the length of CLIP captions and retrain CLIP accordingly. For instance, DCI [40] hired annotators to expand captions based on SAM [16]’s target hints. LaCLIP [9] adopted ChatGPT, Bard, and human rewriting. DreamLIP [45] leveraged ShareCaptioner [5] and InstructBLIP to augment 30M captions. Recap-DataComp-1B [20] used Llama3-trained LLAVA1.5 to extend 1B captions. In order to handle captions longer than 77 tokens, these methods compromised by summarizing long captions into shorter ones [40], splitting captions into multiple segments [9, 45], or finetuning model’s positional encoding to support longer token inputs [44]. LLM2CLIP, in comparison, leverages Llama3 as the text encoder, which enables comprehensive understanding of long and dense captions. It not only resolves the limition of token length but also allows for better understanding using LLMs’ open-world knowledge. As a result, our model achieves a breakthrough in performance.
3 Methods
LLM2CLIP is a fine-tuning strategy for CLIP, consisting of two stages. The first stage aims to improve the separability of the LLM’s output space, and the second stage involves fine-tuning by replacing CLIP’s text model with the LLM.
3.1 Caption Contrastive Fine-tuning
In this section, we aim to fine-tune the LLM’s token outputs to better capture features that can distinguish between image captions.
The process of improving the discriminability of LLM output features on caption text is quite straightforward: we need the distances between captions of the same image to be closer, and those of different images to be further apart. Therefore, we apply a caption contrastive (CC) fine-tuning to the LLM’s output features, treating different captions of the same image as positive samples and the rest of the captions as negative samples.
To obtain enough varied descriptions for the same image, we use the ShareCaptioner [45, 5] modified CC-3M [36] dataset, which provides both original captions and augmented dense captions for each image. These can be treated as positive pairs. We followed the training methodology of LLM2Vec [3], first expanding the LLM’s attention mechanism to bidirectional attention, and employing Masked Next Token Prediction (MNTP) for initialization to achieve better results. Specifically:
  1. 1.
    Bidirectional Attention Transformation: First, we transform the LLM’s causal attention mechanism into bidirectional attention, as we no longer need it to retain generative capabilities but rather function as an encoder. Since autoregressive training is no longer required, switching to bidirectional attention improves its ability to capture contextual information.
  2. 2.
    Masked Next Token Prediction (MNTP): Second, we employ MNTP to train the newly added bidirectional attention mechanism, providing a strong initialization. For a given sequence of $N$ tokens, we mask a subset and predict their values, similar to BERT [7]. However, unlike BERT, we adapt this process to fit the nature of LLMs by predicting tokens just before the masked token. We train on both image captions and pure text with equal weighting. In addition to CC-3M, we also use the Wikitext-103 [28] dataset to preserve the LLM’s text capabilities and prevent divergence from its original strengths.
  3. 3.
    Caption Contrastive Learning: Finally, we perform the actual caption contrastive fine-tuning, using a supervised SimCSE loss to bring captions of the same image closer together and push apart captions of different images. We use two prompt templates: “Given a caption, retrieve a detailed relevant caption” and “Given a detailed caption, retrieve a short relevant caption,” which are prepended to the query (either original or dense caption) to retrieve the corresponding dense or original caption. Similarly, we use a 1.5M-paired general text dataset curated from [37] to maintain strong performance in pure language tasks.
All training is efficiently conducted using LoRA and is completed in just one epoch, ensuring low computational cost. In our experiments, we found that the most critical step was the third step with the SimCSE loss, which significantly enhanced the caption discriminability of the LLM.
3.2 LLM2CLIP: Use LLM to Finetuning
With the modifications to the LLM discussed earlier, we have now obtained a super text encoder that is well-suited for CLIP training. The next step is to use this LLM in conjunction with the pretrained state-of-the-art CLIP visual encoder to reconstruct a more powerful cross-modal feature space.
As shown in  Figure 3, in the LLM2CLIP training phase, we freeze the gradients of the LLM to preserve its inherent capabilities, for two primary reasons. First, this significantly reduces the computational cost and memory footprint of fine-tuning. CLIP training requires a very large batch size to maintain the effectiveness of negative samples. Allocating memory to the LLM could compromise CLIP’s performance. Second, by freezing the LLM, we ensure that the open-world knowledge it has acquired from large-scale corpora remains intact during the multimodal alignment process.
To compensate for the frozen LLM, and inspired by methods like FuseMix [41] and APE [35], we introduce several new linear layers after the LLM as adapters. These layers serve as learnable parameters to improve the alignment between the LLM and CLIP visual encoder. Following the original design of CLIP, we also employ a projector layer to align the dimensions of the two encoders, facilitating the use of CLIP loss for training.
With this powerful LLM-based super text encoder, we achieve a substantial qualitative leap in CLIP’s language comprehension capabilities as results in Table 2. The LLM’s open-world knowledge allows the CLIP visual encoder to learn more structured and globally informed visual representations that are aligned with human knowledge. Moreover, this approach enables us to fully utilize high-quality, long, and dense caption datasets without requiring any special architectural adjustments, which previous works like DCI, DreamLip, and Recaption struggled to effectively leverage.
3.3 Overview and Efficiency Discussion
We propose LLM2CLIP as a method that efficiently incorporates large language models (LLMs) into CLIP training, leveraging the capabilities of LLMs to make cross-modal representation learning significantly more powerful. In our experiments, We evaluated large language models, including Llama with 1B and 8B parameters, as well as Mistral-Nemo with 12B parameters.It might seem that incorporating such large LLMs would greatly increase the computational burden of training CLIP, especially since CLIP itself is computationally expensive, requiring a large batch size. However, our proposed LLM2CLIP is remarkably lightweight. The training overhead isnearly identical to fine-tuning the original CLIP model, with minimal additional cost, yet the LLM provides much stronger supervision.
Here, we highlight some of the design details that significantly improve training efficiency: 1). During the caption contrastive fine-tuning stage, we employ LoRA training for the LLM. Even for a 12B LLM, training with a batch size of 512 only requires around 70GB of GPU memory, making it feasible to run on a single 80GB 8 A100 GPU node. 2). In the LLM2CLIP stage, we freeze the LLM’s gradients and only train the learnable adapter, CLIP’s original Vision encoder, and two projectors. The additional trainable parameters are roughly equivalent to those in the original CLIP, minimizing the overhead. To further reduce the inference cost of using LLMs, we pre-extract all the text features from the training data and store them in disk memory. This ensures that, even though LLMs provide powerful textual supervision, the memory and computational costs during training remain nearly identical to those of standard CLIP training.
For instance, when we trained LLM2CLIP using a Mistral-Nemo [^1] https://mistral.ai/news/mistral-nemo/ 12B model integrated with the commonly used EVA ViT-L/14-224, with a batch size of 4096 on 8 H100 GPUs, the memory usage per GPU was only 30GB, and the entire training process took just 9 hours. Despite this efficient training cost, LLM2CLIP brought transformative improvements in downstream tasks such as long and short text retrieval, cross-language retrieval, and Llava training.
@@ 3.3 Overview and Efficiency Discussion
Methods Flickr30k COCO ShareGPT4V Urban-1k DOCCI Average
@@
ViT-B/16
ALIGN 80.6 62.2 52.0 43.2 75.9 80.6 62.2 59.1 59.7 62.1 66.1 61.4
BLIP 80.6 74.1 61.7 48.5 65.8 74.3 45.5 48.5 50.5 53.5 60.8 59.8
Jina-CLIP 80.6 67.4 55.6 41.1 - - 87.7 88.0 78.7 80.0 75.7 69.1
Long-CLIP 85.8 70.6 56.9 40.9 94.8 93.5 79.1 79.1 63.1 71.4 75.9 71.1
CLIP 82.3 62.2 52.4 33.1 84.5 79.8 67.5 53.1 60.7 57.1 69.5 57.1
+LLM2CLIP 89.2 78.1 62.2 48.7 98.1 97.4 86.1 90.0 84.1 85.0 83.9 79.8
EVA02 86.2 71.5 58.7 42.1 90.5 85.5 67.0 60.8 67.7 68.0 74.0 65.6
+LLM2CLIP 88.5 78.0 63.6 49.8 98.0 98.1 84.7 89.7 85.5 86.8 84.1 80.5
ViT-L/14
Long-CLIP 90.0 76.2 62.8 46.3 97.2 97.3 82.5 86.1 66.5 78.6 79.8 76.9
CLIP 85.2 65.0 56.3 36.5 84.2 83.6 68.3 55.6 63.1 65.8 71.4 61.3
+LLM2CLIP 92.6 81.7 64.9 52.5 98.4 98.4 87.6 92.0 87.6 88.7 86.2 82.7
EVA02 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5
+LLM2CLIP-3M 89.6 77.3 59.7 48.0 98.3 98.6 87.1 91.1 84.9 87.8 83.9 80.6
+LLM2CLIP 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2
+LLM2CLIP-30M 92.0 83.5 69.0 55.3 98.9 98.8 93.1 95.0 89.3 91.2 88.5 84.8
+LLM2CLIP-60M 94.4 83.2 70.4 55.7 99.2 99.4 94.1 95.2 90.2 92.0 89.7 85.1
ViT-L/14-336
CLIP 87.7 67.0 58.0 37.1 86.2 84.0 72.8 57.0 67.4 65.7 74.4 62.2
+LLM2CLIP 91.2 82.1 65.5 53.6 98.1 98.4 90.3 93.2 87.7 89.0 86.6 83.3
+LLM2CLIP-60M 93.9 82.3 68.5 54.8 98.9 99.1 94.6 95.9 89.6 90.6 89.1 84.5
EVA02 89.6 78.0 64.2 47.9 91.5 89.4 76.6 70.0 74.7 76.4 79.3 72.3
+LLM2CLIP 93.9 83.8 68.7 55.7 98.8 99.2 89.5 94.2 89.2 91.3 88.0 84.8
Table 2: Systematic Comparison Experiment Demonstrating the Performance Improvements of LLM2CLIP. Results with a blue background represent the outcomes of LLM2CLIP.
4 EXPERIMENTS
4.1 Experiment Setting
4.1.1 Caption Contrastive Fine-tuning
Training Dataset.
For contrastive learning, we utilized the ShareCaptioner-modified CC-3M dataset [45, 5], which includes both original captions and augmented dense captions for each image. In addition, we used the Wikitext-103 dataset [28] and the E5 dataset [37] during the Masked Next Token Prediction and caption contrastive fine-tuning stages to ensure coverage of the general text domain without excessive bias.
Training Setting.
The language model was trained with LoRA for lightweight training, completing one epoch across all datasets. We used the average of all LLM output tokens as the global text embedding for a caption. All training parameters followed the configuration detailed in [3]. The design of the Adaptor is the same as FuseMix [41], using a simple inverted bottleneck MLP architecture.
4.1.2 LLM2CLIP Fine-tuning
Training Dataset.
We defined three experimental configurations based on dataset size:
  • LLM2CLIP-3M: This configuration uses only the CC-3M dataset, serving as our lightweight setting.
  • LLM2CLIP-15M: Our default configuration, combining CC-3M and CC-12M datasets.
  • LLM2CLIP-60M: A larger configuration that combines CC-3M, CC-12M, YFCC-15M, and a 30M randomly selected subset from Recaption-1B [20].
All datasets included dense captions rewritten by multimodal large language models (MLLMs). Specifically, the CC-3M, CC-12M, and YFCC datasets were sourced from [45] and rewritten using ShareCaptioner, while Recaption data were rewritten with Llama3-LLAVA1.5. A 1:1 mix of original captions and dense captions was used by default.
Training Setting.
Each configuration was trained for four epochs. We froze the LLM gradients to reduce computational overhead and pre-extracted text features using the LLM, storing them on disk to avoid repeated inference during training. Inspired by Fusemix’s efficient training strategy, we added a four-layer linear adapter after the LLM and trained with gradients enabled for the CLIP ViT model to better integrate LLM knowledge. Additionally, we introduced two projector layers after each encoder for dimensional alignment.
Training the CLIP ViT-L/14-224 model required approximately 30GB of GPU memory per GPU and took 2, 9, and 45 hours on 8 H100 GPUs for the 3M, 15M, and 60M datasets, respectively. We employed the AdamW [26] optimizer, utilizing a total batch size of 4096, and set the learning rate to 1e-5 with a cosine decay rate of 0.05.
For language models, we tested four configurations: Jina-Embeddings-V2, and three popular models from the Llama3 family [8]: Llama 3.2 1B, Llama 3 8B, and Mistral-Nemo 12B. Unless otherwise noted, experiments used EVA02 ViT-L/14 as the baseline and Llama 3 8B as the default language model.
Methods Flickr-CN COCO-CN
I2T@1 I2T@5 I2T@10 T2I@1 T2I@5 T2I@10 I2T@1 I2T@5 I2T@10 T2I@1 T2I@5 T2I@10
ViT-L/14-336
Wukong 76.1 94.8 97.5 51.7 78.9 86.3 53.4 80.2 90.1 55.2 81.0 90.6
CN-CLIP 80.2 96.6 98.2 68.0 90.7 95.4 63.4 84.2 92.9 64.0 89.2 94.4
JinaCLIP 3.30 9.90 15.1 0.7 3.5 6.0 2.9 8.9 13.7 1.0 4.9 8.2
EVA02 4.40 11.8 16.7 0.94 2.9 4.8 2.7 9.8 15.2 1.0 3.7 7.3
+LLM2CLIP 86.9 98.1 99.3 75.1 92.9 96.0 69.1 92.5 97.2 70.0 92.6 96.7
Table 3: Retrieval Performance across Flickr30K-CN and COCO-CN.
@@ 4.1.2 LLM2CLIP Fine-tuning
Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average CRA

### v4

@@ Abstract
CLIP is a foundational multimodal model that aligns image and text features into a shared representation space via contrastive learning on large-scale image-text pairs. Its effectiveness primarily stems from the use of natural language as rich supervision. Motivated by the remarkable advancements in large language models (LLMs), this work explores how LLMs’ superior text understanding and extensive open-world knowledge can enhance CLIP’s capability, especially for processing longer and more complex image captions. We propose an efficient post-training strategy that integrates LLMs into pretrained CLIP. To address the challenge posed by the autoregressive nature of LLMs, we introduce a caption-to-caption contrastive fine-tuning framework, significantly enhancing the discriminative quality of LLM outputs. Extensive experiments demonstrate that our approach outperforms LoRA-based methods, achieving nearly fourfold faster training with superior performance. Furthermore, we validate substantial improvements over state-of-the-art models such as CLIP, EVA02, and SigLip2 across various zero-shot multimodal retrieval tasks, cross-lingual retrieval tasks, and multimodal language model pretraining.
@@ 1 Introduction
CLIP [28] is one of the most influential multimodal foundational models today. What makes CLIP particularly powerful as a vision encoder? The core of its strength lies in its unprecedented ability to align visual pretraining with natural language—the carrier of human knowledge. Unlike earlier vision encoders such as VGG and ResNet, which relied on the limited ImageNet dataset and simple image categories with just a few words, CLIP is trained on web-scale data using rich descriptive text. This alignment with language is what sets CLIP apart and unlocks its vast potential. However, since CLIP’s introduction, large language models (LLMs) have advanced significantly. Models like GPT-4 [1] and Llama [6] now demonstrate remarkable language capabilities, yet these advancements have not translated into corresponding improvements in visual representation learning. This raises the question: Can the capabilities of LLMs be harnessed to further improve multimodal representation learning?
@@
This paper successfully explores a efficient method of leveraging LLM capabilities in CLIP pretraining. LLMs, as text generation models, are primarily pretrained to capture the distribution of natural language, thus their outputs naturally lie close to textual spaces. However, what CLIP pretraining actually requires is a robust text embedding model. Previous works such as LLM2Vec [2] and NV-Embed-v2 [17] have demonstrated that fine-tuning LLMs significantly enhances their ability as text encoders. Nevertheless, it remains unclear how directly applying LLMs affects CLIP performance. Our preliminary results indicate that directly integrating LLMs into CLIP pretraining can be detrimental because LLMs struggle to differentiate effectively between image descriptions, as illustrated by example cases in Figure 3. Moreover, as shown in Table 1, we conducted caption-to-caption retrieval experiments on the COCO 5K test set (each image has five captions; we used the first two captions as positive pairs) and observed that large language models like Llama 3 8B achieve very low accuracy, barely distinguishing between different captions of the same image.
Language Model Top 1 Acc.
@@
Table 1: Comparison of top-1 retrieval accuracy for various language models on the MS COCO [22] 5K test set.
@@
We designed a caption contractive (CC) fine-tuning strategy, applying lightweight fine-tuning to the output tokens of LLM using image captioning dataset as in  Figure 2. The primary goal of this training task was to adjust the output space, improving the model’s ability to distinguish between different captions. We utilized a supervised SimCSE [8, 2] contrastive learning loss, where the original captions and re-annotated captions generated by ShareCaptioner [3] were treated as positive pairs, pulling them closer. In contrast, all other captions formed a negative sample set that the model learned to push away. Remarkably, after this CC fine-tuning, the caption retrieval accuracy, as shown in Table 1, rose from 5.2% to 29.5%. This successful fine-tuning process enables us to more effectively harness the capabilities of LLMs for CLIP training.
Building on fine-tuning of LLMs, we designed an efficient approach, LLM2CLIP, to enhance the performance of pretrained CLIP. As in  Figure 1, This method uses an LLM to further fine-tune the pretrained CLIP, allowing the LLM to impart knowledge and capabilities to CLIP’s vision encoder that were lacking during its initial pretraining. LLM2CLIP consists of two stages: in the first stage, we apply caption contrastive (CC) fine-tuning on the LLM with a small dataset to adjust the LLM’s output space. In the second stage, we fine-tune the CLIP vision encoder with the CC-fine-tuned LLM, with LLM gradients frozen, to further improve CLIP’s performance.
@@
    We successfully verified that LLM2CLIP can substantially enhance the performance of CLIP and EVA02 across multiple tasks, including short-text retrieval, long-text retrieval, cross-language retrieval, and multimodal large language model pretraining, through a simple post-training strategy. Remarkably, LLM2CLIP can further improve the short-text retrieval accuracy of state-of-the-art large-scale pretrained models like SigLip2 [33] by +1 and +1.9 points, achieve impressive gains of +14.8 and +15.8 points in long-text retrieval, and deliver significant improvements of +11.9 and +15.2 points in multilingual tasks.
@@ 2 Related Works
CLIP meets Stronger Language Models. Several works have explored the integration of LLMs into CLIP. JinaCLIP [14] employed Jina-embeddings-v2 [9] as the text encoder, which is a BERT variant with 137M parameters, supporting longer texts. Though achieving similar visual performance to EVA-CLIP [30], its text encoder is far behind our used Llama3-8B, limiting the potential benefits from LLMs. E5-V [12] took a different approach by aggregating LLM features at the MLLM layer of LLAVA-next-8B [18] while freezing the ViT gradient to focus on adjusting the MLLM output. However, freezing visual encoder does not address the problem that ViT inherently lack complex visual feature extraction capabilities, leading to much worse performance than LLM2CLIP. MATE [10] designed a learnable adaptor to bridge the gap between CLIP’s text encoder and LLMs. They trained the CLIP visual encoder using LoRA on a small dataset focused on long-text image retrieval tasks. However, they did not recognize the critical issue we propose in this paper: the poor separability of LLM feature space, which is insufficient for direct support of CLIP training. Our work aims to thoroughly explore the impact of LLMs on CLIP as a foundational model, going beyond the influence on long-text retrieval.
@@
CLIP meets Longer Captions. It has been widely recognized that the quality of CLIP’s text embedding is coarse and limited to only 77 tokens. Many works have attempted to extend the length of CLIP captions and retrain CLIP accordingly. For instance, DCI [34] hired annotators to expand captions based on SAM [13]’s target hints. LaCLIP [7] adopted ChatGPT, Bard, and human rewriting. DreamLIP [41] leveraged ShareCaptioner [3] and InstructBLIP to augment 30M captions. Recap-DataComp-1B [21] used Llama3-trained LLAVA1.5 to extend 1B captions. In order to handle captions longer than 77 tokens, these methods compromised by summarizing long captions into shorter ones [34], splitting captions into multiple segments [7, 41], or finetuning model’s positional encoding to support longer token inputs [39]. LLM2CLIP, in comparison, leverages Llama3 as the text encoder, which enables comprehensive understanding of long and dense captions. It not only resolves the limition of token length but also allows for better understanding using LLMs’ open-world knowledge. As a result, our model achieves a breakthrough in performance.
@@
Methods Resolution Flickr30k COCO ShareGPT4V Urban-1k DOCCI Average
@@
ViT-B/16(86M)
ALIGN [11] 224 80.6 62.2 52.0 43.2 75.9 80.6 62.2 59.1 59.7 62.1 66.1 61.4
BLIP [19] 224 80.6 74.1 61.7 48.5 65.8 74.3 45.5 48.5 50.5 53.5 60.8 59.8
Long-CLIP [39] 224 85.8 70.6 56.9 40.9 94.8 93.5 79.1 79.1 63.1 71.4 75.9 71.1
jina-clip-v2 [15] 224 84.4 69.8 57.1 41.9 94.5 91.1 80.4 78.0 77.6 78.2 78.8 71.8
SigLIP [38] 224 89.1 74.7 65.8 47.8 85.8 83.5 62.8 62.4 70.1 70.6 74.7 67.8
MetaCLIP [36] 224 85.6 70.8 59.3 41.3 90.4 86.6 68.9 63.3 70.9 71.5 75.0 66.7
CLIP [28] 224 82.3 62.2 52.4 33.1 84.5 79.8 67.5 53.1 60.7 57.1 69.5 57.1
     +LLM2CLIP-15M 224 91.8 80.7 64.8 52.4 97.8 97.6 90.1 92.1 86.1 86.9 86.1 81.9
EVA02 [31] 224 86.2 71.5 58.7 42.1 90.5 85.5 67.0 60.8 67.7 68.0 74.0 65.6
     +LLM2CLIP-15M 224 91.2 81.4 66.0 53.9 98.5 98.5 88.2 91.3 86.8 88.5 86.1 82.7
ViT-L/14 (307M)
Long-CLIP [39] 224 90.0 76.2 62.8 46.3 97.2 97.3 82.5 86.1 66.5 78.6 79.8 76.9
MetaCLIP [36] 224 90.1 76.4 64.5 47.1 86.4 79.5 73.4 70.0 76.5 76.7 78.2 69.9
SAIL-L-NV2 [40] 224 87.9 75.8 62.4 48.7 94.8 94.9 81.5 80.2 76.5 78.9 80.6 75.7
CLIP [28] 224 85.2 65.0 56.3 36.5 84.2 83.6 68.3 55.6 63.1 65.8 71.4 61.3
     +LLM2CLIP-15M 224 93.0 83.5 67.4 56.4 98.7 98.6 91.5 94.1 88.3 90.4 87.8 84.6
EVA02 [31] 224 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5
     +LLM2CLIP-3M 224 94.3 84.3 65.2 56.0 98.0 98.0 91.0 94.7 87.8 90.7 87.3 84.7
     +LLM2CLIP-15M 224 94.6 85.0 69.5 58.3 98.9 99.1 93.6 95.7 89.8 91.2 89.3 85.8
     +LLM2CLIP-60M 224 95.9 85.1 71.7 58.5 99.2 99.2 95.2 97.0 90.1 92.0 90.4 86.3
CLIP [28] 336 87.7 67.0 58.0 37.1 86.2 84.0 72.8 57.0 67.4 65.7 74.4 62.2
     +LLM2CLIP-15M 336 91.2 82.1 65.5 53.6 98.1 98.4 90.3 93.2 87.7 89.0 86.6 83.3
     +LLM2CLIP-60M 336 93.9 84.3 68.5 54.8 98.9 99.1 96.6 97.9 89.6 91.6 91.1 84.5
EVA02 [31] 336 89.6 78.0 64.2 47.9 91.5 89.4 76.6 70.0 74.7 76.4 79.3 72.3
     +LLM2CLIP-15M 336 94.7 85.1 69.7 58.5 98.8 99.2 93.5 96.2 89.2 91.3 88.0 84.8
     +LLM2CLIP-60M 336 95.9 85.4 72.4 58.8 98.8 99.2 95.5 97.3 90.7 92.6 91.0 86.8
SigLIP [38] 384 93.7 81.4 72.0 53.9 90.1 90.4 76.3 74.4 77.8 79.5 82.0 75.9
SO/14(428M)
SigLIP [38] 224 91.0 75.2 69.8 51.8 56.1 50.0 31.5 29.1 42.3 44.8 58.1 50.2
SigLIP2 [33] 224 93.9 82.9 72.0 55.5 90.2 87.2 75.7 74.5 77.0 78.9 81.8 75.8
     +LLM2CLIP-60M 224 95.2 84.6 73.7 57.6 99.4 99.4 96.2 97.3 91.5 92.9 91.2 86.4
SigLIP [38] 384 94.3 83.0 72.4 54.3 91.6 89.4 74.5 73.6 74.5 75.8 81.5 75.2
InternVL(6B)
InternVL [4] 224 94.2 81.6 69.4 53.2 92.4 90.5 79.6 77.7 79.2 81.6 83.0 76.9
Table 2: Systematic comparison of model performance on multiple datasets. The light blue row indicate LLM2CLIP post trained version.
Models Flickr-CN [16] COCO-CN [20] XM3600 [32]
I2T T2I I2T T2I I2T T2I
CN-CLIP 80.2 68.0 63.4 64.0
EVA-L-224 4.4 0.9 2.6 1.0 14.0 8.0
     +LLM2CLIP 90.6 75.6 72.0 70.1 68.3 56.0
Siglip2 79.2 56.9 55.3 51.7 59.7 48.2
     +LLM2CLIP 90.0 76.1 70.8 70.2 69.1 56.3
Table 3: Multi-lingua retrieval results on different datasets.
@@
Methods Flickr COCO ShareGPT4V Urban-1k DOCCI Average